OpenStudy (superhelp101):

Can someone please help me?? Events M and N have probabilities such that P(M)=.4, P(N)=.28. P(M U N)=.56, and P(M n N)=.12 Are events M and N independent? A No, because P(M)-P(N)=P(M n N) B no, because P(M) times P(N) is not equal to P(M n N) C yes because P(M)+P(N)= P(M U N) D yes because P(M) times P(N) is not equal to P(M U N)

OpenStudy (superhelp101):

Anyone I am stuck any help?

OpenStudy (anonymous):

they are independent if \[\frac{P(M\cap N)}{P(N)}=P(M)\]

OpenStudy (anonymous):

or if \[P(M)P(N)=P(M\cap N)\]

OpenStudy (superhelp101):

ohh okay I see let me agree on a answer and I will check with you to verify

OpenStudy (superhelp101):

my problem has \[\neq \] in it

OpenStudy (superhelp101):

not and equal sign tho

OpenStudy (anonymous):

yes, they are not equal, so they are dependent

OpenStudy (superhelp101):

okay so the second choice would make sense am i right